# Controllability of the Fisher-Stefan system

### v1

@@ 1.2 Main result and reformulation of the problem
For that, we apply the Lyusternik-Graves Inverse Mapping Theorem (see Theorem 3.1).
Structure of this paper. The rest of the paper is structured as follows. Section 2 focuses on the well-posedness of the linearized system of (1.11), as well as the adjoint system. We also prove the null controllability of the linearized system of (1.11) with a source term. In Section 3, we show the local null controllability of the nonlinear system (1.11). Section 4 focuses on numerical analysis, introducing recent deep learning techniques, especially Physics-Informed Neural Networks (PINNs), to simulate the behavior of the density and free boundary in the nonlinear system, as well as its linear approximation. Finally, Section 5 is devoted to the conclusion.
2 Study of the linearized system of (1.11)
As usual, the local inversion of a nonlinear mapping relies on the surjectivity (which is sufficient) of its derivative around the point of interest. At this stage, the surjectivity of our mapping defined in (1.13) is connected to the null controllability of the inhomogeneous linearized system (around zero) of (1.11).
@@ 2.1 Linearization
The inhomogenuous linearized system at $(0,0)$ of (1.11) is given by:
$&\overline{h}z_{t}-z_{xx}-\frac{x}{2}(\overline{h})^{\prime}z_{x}+\mu x\overline{\psi}_{x}z_{x}(1,\cdot)+mk+nz=f_{0}+\mathds{1}_{\omega}v&&\text{in}\;\tilde{Q},\\ &z(-1,\cdot)=z(1,\cdot)=0&&\text{in}\;(0,T),\\ &k^{\prime}+2\mu z_{x}(1,\cdot)=g_{0}&&\text{in}\;(0,T),\\ &k(0)=k_{0},\\ &z(\cdot,0)=z_{0}&&\text{in}\;(-1,1),\end{aligned}\right.$ (2.1)
where $m$ and $n$ are defined in (1.9), and the functions $f_{0}$ and $g_{0}$ belong to suitable function spaces that exhibit exponential decay as $t$ approaches $T$ from the left, as will be defined more precisely later. In the sequel, we use the following notations:
Notation.
@@
$\textbf{L}_{1}(z,k)$ $:=\overline{h}z_{t}-z_{xx}-\frac{x}{2}(\overline{h})^{\prime}z_{x}+\mu x\overline{\psi}_{x}z_{x}(1,\cdot)+mk+nz,$ (2.2)
$\textbf{L}_{2}(z,k)$ $:=k^{\prime}+2\mu z_{x}(1,\cdot),$
$\textbf{L}_{1}^{\star}(z,k)$ $:=-(\overline{h}z)_{t}-z_{xx}+\frac{(\overline{h})^{\prime}}{2}(xz)_{x}+nz,$
$\textbf{L}_{2}^{\star}(z,k)$ $:=-k^{\prime}+\int_{-1}^{1}m(x,\cdot)z(x,\cdot)\mathrm{d}x.$
We are focused on addressing the following categories of solutions for (2.1).
Definition 2.1.
@@
    A solution by transposition of (2.1) is a couple of functions $(z,k)\in L^{2}(\tilde{Q})\times L^{2}(0,T)$ such that for any $(w,\theta)\in H^{2,1}_{0}(\tilde{Q})\times H^{1}(0,T)$ with $w(\cdot,T)=\theta(T)=0$, $w(-1,t)=0$ and $w(1,t)-\mu\displaystyle\int_{-1}^{1}x\overline{\psi}_{x}w(x,t)\mathrm{d}x-2\mu\theta(t)=0$, $t\in(0,T)$, we have
@@
    A strong solution of (2.1) is a couple of functions $(z,k)\in H^{2,1}_{0}(\tilde{Q})\times H^{1}(0,T)$ fulfilling (2.1)1italic-(2.1subscriptitalic-)$_{1}$ in $L^{2}(\tilde{Q})$, (2.1)3italic-(2.1subscriptitalic-)$_{3}$ in $L^{2}(0,T)$ and initial conditions.
@@
A transposition solution is related to the solutions of adjoint system, where conditions $w(-1,t)=0$ and $w(1,t)-\mu\displaystyle\int_{-1}^{1}x\overline{\psi}_{x}w(x,t)\mathrm{d}x-2\mu\theta(t)=0$, $t\in(0,T)$ appear as boundary conditions of the adjoint system (see system (2.4)).
System (2.1) has a unique strong solution in $H_{0}^{2,1}(\tilde{Q})\times H^{1}(0,T)$, for any $f_{0}\in L^{2}(\tilde{Q}),v\in L^{2}(\omega\times(0,T)),g_{0}\in L^{2}(0,T),k_{0}\in\mathbb{R}$ and any $z_{0}\in H^{1}_{0}(-1,1)$, see Proposition 2.7 in [21]. Moreover, there is a positive constant $C$ depending on $\overline{\psi},\overline{h},\mu$ and $T$ such that
$\|z\|_{H_{0}^{2,1}(\tilde{Q})}+\|k\|_{H^{1}(0,T)}\leq C\left(\|f_{0}\|_{L^{2}(\tilde{Q})}+\|v\|_{L^{2}(\omega\times(0,T))}+\|g_{0}\|_{L^{2}(0,T)}+|k_{0}|+\|z_{0}\|_{H^{1}_{0}(-1,1)}\right).$ (2.3)
2.2 Adjoint system of (2.1)
To determine the adjoint system of (2.1), multiply the first equation of (2.1) by $w:=w(x,t)$ and the third equation by $\theta:=\theta(t)$ and proceeding with integrations by parts, we find that the adjoint system of (2.1) is given by
{$&-(\overline{h}w)_{t}-w_{xx}+\frac{(\overline{h})^{\prime}}{2}(xw)_{x}+nw=f_{1}&&\text{in}\;\tilde{Q},\\ &w(-1,t)=0&&\text{in}\;(0,T),\\ &w(1,t)=\mu\int_{-1}^{1}x\overline{\psi}_{x}(x,t)w(x,t)\mathrm{d}x+2\mu\theta(t)&&\text{in}\;(0,T),\\ &-\theta^{\prime}(t)+\int_{-1}^{1}m(x,t)w(x,t)\mathrm{d}x=g_{1}&&\text{in}\;(0,T),\\ &\theta(T)=\theta_{T},\\ &w(\cdot,T)=w_{T},&&\text{in}\;(-1,1),\end{aligned}\right.$ (2.4)
@@
$w_{T}(-1)=0\quad\mbox{and}\quad w_{T}(1)=\mu\int_{-1}^{1}x\overline{\psi}_{x}(x,T)w_{T}(x)\mathrm{d}x+2\mu\theta_{T}.$ (2.5)
The following result assures the existence of strong solutions of (2.4) and its proof is a direct application of [21, Proposition 2.5] based on Leray-Schauder’s Fixed-Point Principle.
Proposition 2.2.
Let $w_{T}\in H^{1}(-1,1)$ and $\theta_{T}\in\mathbb{R}$ such that the compatibility condition (2.5) is satisfied and $(f_{1},g_{1})\in L^{2}(\tilde{Q})\times L^{2}(0,T)$, then (2.4) has a unique strong solution $(w,\theta)\in H^{2,1}(\tilde{Q})\times H^{1}(0,T)$, which satisfies the following estimate:
$\|w\|_{H^{2,1}(\tilde{Q})}+\|\theta\|_{H^{1}(0,T)}\leq C\left(\|f_{1}\|_{L^{2}(\tilde{Q})}+\|g_{1}\|_{L^{2}(0,T)}+|\theta_{T}|+\|w_{T}\|_{H^{1}(-1,1)}\right).$ (2.6)
@@
2.3 Null controllability of system (2.1)
The objective of this section is to prove null controllability for system (2.1) and some state estimation and control regularity that requires certain regularity of the initial data and inhomogeneous terms. We consider the following positive weight functions $\alpha$ and $\xi$, which depend on $\omega$:
@@
$\eta>0\text{ in }[-1,1],\quad\min_{x\in[-1,1]\backslash\omega^{\prime}}\left|\eta_{x}(x)\right|>0,\quad\eta(-1)=\eta(1)=\min_{x\in[-1,1]}\eta(x),$ (2.7)
even if we divide by $||\eta||_{\infty}$, we can also assume that $||\eta||_{\infty}=1$, where $\omega^{\prime}\subset\subset\omega$ is an open non empty subset. The existence of such a function $\eta$ satisfying (2.7) is proved in [26].
Notation.
@@
Proposition 2.3.
There exist constants $\lambda_{1},s_{1}\geq 1$ and $C_{1}>0$ such that, for any $\lambda\geq\lambda_{1}$, any $s\geq s_{1}(T+T^{2})$, any $w_{T}\in H^{1}(-1,1)$ and any $\theta_{T}\in\mathbb{R}$ with compatibility condition (2.5) is satisfied and any source terms $(f_{1},g_{1})\in L^{2}(\tilde{Q})\times L^{2}(0,T)$, the strong solution to (2.4) satisfies:
@@
$\leq C_{1}\left(s^{3}\lambda^{4}\iint_{\omega\times(0,T)}e^{-2s\alpha}\xi^{3}|w|^{2}\mathrm{d}x\mathrm{d}t+\iint_{\tilde{Q}}e^{-2s\alpha}|f_{1}|^{2}\mathrm{d}x\mathrm{d}t+\int_{0}^{T}e^{-2s\hat{\alpha}}|g_{1}|^{2}\mathrm{d}t\right).$ (2.8)
@@
System (2.4) can be written in the form
$&\varphi_{t}+d(t)\varphi_{xx}=f&&\text{in}\;\tilde{Q},\\ &\varphi(-1,t)=0&&\text{in}\;(0,T),\\ &\varphi(1,t)=\gamma(t)+\int_{-1}^{1}N(x,\cdot)\varphi(x,\cdot)\mathrm{d}x&&\text{in}\;(0,T),\\ &\gamma^{\prime}(t)-\int_{-1}^{1}R(x,\cdot)\varphi(x,\cdot)\mathrm{d}x=g&&\text{in}\;(0,T),\\ &\gamma(T)=2\mu\theta_{T},\\ &\varphi(\cdot,T)=\varphi_{T}&&\text{in}\;(-1,1),\end{aligned}\right.$ (2.9)
where $(\varphi,\gamma):=(w,2\mu\theta)$, $d:=(\overline{h})^{-1},\;N:=\mu x\overline{\psi}_{x},\;R:=2\mu m$, $f:=-d\left(f_{1}+\left(\frac{(\overline{h})^{\prime}}{2}-n\right)\varphi-\frac{x}{2}(\overline{h})^{\prime}\varphi_{x}\right)$ and $g:=-2\mu g_{1}$. Since $(\overline{\psi},\overline{h})$ is a smooth data (see (1.12)), and the Sobolev continuous embedding $H^{2,1}(\tilde{Q})\hookrightarrow C([0,T];H^{1}(-1,1))$ and $H^{1}(-1,1)$ is a Banach algebra, we obtain
@@
Note that the regularity $W^{1,\infty}(0,T)$ is sufficient for $d$. Hence, we are in a position to apply Theorem 3.1 of [21], so there exist constants $\lambda_{1},s_{1}\geq 1$ and $C>0$ such that, for any $\lambda\geq\lambda_{1}$, any $s\geq s_{1}(T+T^{2})$, the strong solution to (2.9) satisfies:
@@
We will deduce a Carleman estimate with functions blowing up only at $t=T$, which allows us to prove an observability inequality for system (2.4). Define the new weight functions:
@@
Note that $\ell\in C^{1}([0,T])$. An estimate with such weights is given in the following result. In the proof, we will use Proposition 2.3 and energy estimates (2.6).
Proposition 2.4.
There exist constants $\lambda_{1},s_{1}\geq 1$ such that for any $s\geq s_{1}$, any $\lambda\geq\lambda_{1}$, there exists a constant $C:=C(s,\lambda,T)$, which satisfies the following property: for any strong solution $(w,\theta)$ of (2.4), we have the following estimate:
@@
$\leq C\left(\iint_{\omega\times(0,T)}e^{-2s\beta}\ell^{-3}|w|^{2}\mathrm{d}x\mathrm{d}t+\iint_{\tilde{Q}}e^{-2s\beta}|f_{1}|^{2}\mathrm{d}x\mathrm{d}t+\int_{0}^{T}e^{-2s\hat{\beta}}|g_{1}|^{2}\mathrm{d}t\right).$ (2.10)
with $s_{1}$ and $\lambda_{1}$ as in Proposition 2.3 and $\hat{\beta}=\displaystyle\max_{x\in[-1,1]}\beta(x,\cdot)$.
@@
To do so, we split the left-hand side of (2.10) into two parts, corresponding to the restrictions of $(w,\theta)$ on $[-1,1]\times\left(0,T/2\right)$ and $[-1,1]\times\left(T/2,T\right)$.
Firstly, on $[-1,1]\times\left(T/2,T\right)$. Using Carleman’s estimate (2.8), $\alpha=\beta$ and $\xi=e^{\lambda(\kappa+\eta(x))}\ell^{-1}$ in $[-1,1]\times\left(T/2,T\right)$, we obtain
@@
$\leq C\left(\iint_{\omega\times(0,T)}e^{-2s\alpha}\xi^{3}|w|^{2}\mathrm{d}x\mathrm{d}t+\iint_{\tilde{Q}}e^{-2s\alpha}|f_{1}|^{2}\mathrm{d}x\mathrm{d}t+\int_{0}^{T}e^{-2s\hat{\alpha}}|g_{1}|^{2}\mathrm{d}t\right).$ (2.11)
@@
$\alpha\geq\beta\quad\mbox{in}\;\;[-1,1]\times(0,T)\quad\mbox{and}\quad\hat{\alpha}\geq\hat{\beta}\quad\mbox{in}\;\;(0,T).$ (2.12)
@@
$e^{-2s\alpha}\xi^{3}\leq Ce^{-2s\beta}\ell^{-3}\quad\mbox{in}\;\;[-1,1]\times(0,T),$ (2.13)
where $C>0$ only depending on $s$, $\lambda$ and $T$. Considering (2.11)-(2.13), we obtain
@@
Secondly, on $[-1,1]\times\left(0,T/2\right)$. Let us define the function $\vartheta\in C^{1}([0,T])$ such that
@@
{$&-(\overline{h}w^{\vartheta})_{t}-w^{\vartheta}_{xx}+\frac{(\overline{h})^{\prime}}{2}(xw^{\vartheta})_{x}+nw^{\vartheta}=f^{\vartheta}&&\text{in}\;\tilde{Q},\\ &w^{\vartheta}(-1,t)=0&&\text{in}\;(0,T),\\ &w^{\vartheta}(1,t)=\mu\int_{-1}^{1}x\overline{\psi}_{x}(x,t)w^{\vartheta}(x,t)\mathrm{d}x+2\mu\theta^{\vartheta}(t)&&\text{in}\;(0,T),\\ &-\theta^{\vartheta}_{t}(t)+\int_{-1}^{1}m(x,t)w^{\vartheta}(x,t)\mathrm{d}x=g^{\vartheta}&&\text{in}\;(0,T),\\ &\theta^{\vartheta}(T)=0,\\ &w^{\vartheta}(\cdot,T)=0&&\text{in}\;(-1,1),\end{aligned}\right.$
where $(f^{\vartheta},g^{\vartheta}):=\vartheta(f_{1},g_{1})-\vartheta_{t}(\overline{h}w,\theta)\in L^{2}(\tilde{Q})\times L^{2}(0,T)$. From energy estimate (2.6), one has
@@
$\quad\left.+\|g_{1}\|_{L^{2}(0,3T/4)}+\|w\|_{L^{2}((-1,1)\times(T/2,3T/4))}+\|\theta\|_{L^{2}(T/2,3T/4)}\right).$ (2.14)
Using the boundedness from above and from below of the weight functions $\beta$, $\hat{\beta}$ and $\ell$ in $[-1,1]\times\left(0,3T/4\right)$ and (2.14), we obtain
@@
which, combined with (2.14), yields (2.10). ∎
@@
Proposition 2.5.
Let $(z_{0},k_{0})\in H^{1}_{0}(-1,1)\times\mathbb{R}$ and $(f_{0},g_{0})$ are source terms such that $\rho(f_{0},g_{0})\in L^{2}(\tilde{Q})\times L^{2}(0,T)$. Then, there exists a control $v^{*}$ such that the strong solution $(z^{*},k^{*})$ of (2.1) corresponding to $v^{*}$, $(f_{0},g_{0})$ and $(z_{0},k_{0})$, satisfies:
@@
$\quad\quad\leq C\left(\|\rho f_{0}\|_{L^{2}(\tilde{Q})}^{2}+\|\rho g_{0}\|_{L^{2}(0,T)}^{2}+\|z_{0}\|_{L^{2}(0,T)}^{2}+|k_{0}|^{2}\right).$ (2.15)
In particular (2.1) is null controllable. Moreover, $\rho_{2}(z^{*},k^{*})\in H^{2,1}(\tilde{Q})\times H^{1}(0,T)$
$\|\rho_{2}z^{*}\|^{2}_{H^{2,1}(\tilde{Q})}+\|\rho_{2}k^{*}\|^{2}_{H^{1}(0,T)}\leq C\left(\|\rho f_{0}\|_{L^{2}(\tilde{Q})}^{2}+\|\rho g_{0}\|_{L^{2}(0,T)}^{2}+\|z_{0}\|_{L^{2}(0,T)}^{2}+|k_{0}|^{2}\right).$ (2.16)

### v2

@@ 1.1 Main result and reformulation of the problem
For that, we apply the Lyusternik-Graves Inverse Mapping Theorem (see Theorem 7). This approach to proving local null controllability has been used in several works by Fernández-Cara and co-authors, notably for quasilinear heat equations and the one-phase Stefan problem [1, 19]. In the context of local null controllability for volume-surface reaction-diffusion equations with dynamic boundary conditions, we refer to [12].
Structure of this paper. The rest of the paper is structured as follows. Section 2 focuses on the well-posedness of the linearized system of (11), as well as the adjoint system. We also prove the null controllability of the linearized system of (11) with a source term. In Section 3, we show the local null controllability of the nonlinear system (11). Section 4 focuses on numerical analysis, introducing recent deep learning techniques, especially Physics-Informed Neural Networks (PINNs), to simulate the behavior of the density and free boundary in the nonlinear system, as well as its linear approximation. Finally, Section 5 is devoted to the conclusion.
2 Study of the linearized system of (11)
As usual, the local inversion of a nonlinear mapping relies on the surjectivity (which is sufficient) of its derivative around the point of interest. At this stage, the surjectivity of our mapping defined in (13) is connected to the null controllability of the inhomogeneous linearized system (around zero) of (11).
@@ 2.1 Linearization
The inhomogenuous linearized system at $(0,0)$ of (11) is given by:
$\left\{\begin{aligned} &\overline{h}z_{t}-z_{xx}-\frac{x}{2}(\overline{h})^{\prime}z_{x}+\mu x\overline{\psi}_{x}z_{x}(1,t)+mk+nz=f_{0}+\mathds{1}_{\omega}v&&\text{in}\;\tilde{Q},\\ &z(-1,t)=z(1,t)=0&&\text{in}\;(0,T),\\ &k^{\prime}(t)+2\mu z_{x}(1,t)=g_{0}&&\text{in}\;(0,T),\\ &k(0)=k_{0},\\ &z(\cdot,0)=z_{0}&&\text{in}\;(-1,1),\end{aligned}\right$. (14)
where $m$ and $n$ are defined in (9), and the functions $f_{0}$ and $g_{0}$ belong to suitable function spaces that exhibit exponential decay as $t$ approaches $T$ from the left, as will be defined more precisely later. In the sequel, we use the following notations:
Notation 1.
@@
$\textbf{L}_{1}(z,k)$ $=\overline{h}z_{t}-z_{xx}-\frac{x}{2}(\overline{h})^{\prime}z_{x}+\mu x\overline{\psi}_{x}z_{x}(1,\cdot)+mk+nz,$ (15)
$\textbf{L}_{2}(z,k)$ $=k^{\prime}+2\mu z_{x}(1,\cdot),$
$\textbf{L}_{1}^{\star}(z,k)$ $=-(\overline{h}z)_{t}-z_{xx}+\frac{(\overline{h})^{\prime}}{2}(xz)_{x}+nz,$
$\textbf{L}_{2}^{\star}(z,k)$ $=-k^{\prime}+\int_{-1}^{1}m(x,\cdot)z(x,\cdot)\mathrm{d}x.$
We are focused on addressing the following categories of solutions for (14).
Definition 2.
@@
    A solution by transposition of (14) is a couple of functions $(z,k)\in L^{2}(\tilde{Q})\times L^{2}(0,T)$ such that for any $(w,\theta)\in H^{2,1}_{0}(\tilde{Q})\times H^{1}(0,T)$ with $w(\cdot,T)=\theta(T)=0$, w(1,t)=0w(-1,t)=0 and $w(1,t)-\mu\displaystyle\int_{-1}^{1}x\overline{\psi}_{x}w(x,t)\mathrm{d}x-2\mu\theta(t)=0$, $t\in(0,T)$, we have
@@
[…]
@@
A transposition solution is related to the solutions of adjoint system, where conditions w(1,t)=0w(-1,t)=0 and $w(1,t)-\mu\displaystyle\int_{-1}^{1}x\overline{\psi}_{x}w(x,t)\mathrm{d}x-2\mu\theta(t)=0$, $t\in(0,T)$ appear as boundary conditions of the adjoint system (see system (17)).
System (14) has a unique strong solution in $H_{0}^{2,1}(\tilde{Q})\times H^{1}(0,T)$, for any $f_{0}\in L^{2}(\tilde{Q}),v\in L^{2}(\omega\times(0,T)),g_{0}\in L^{2}(0,T),k_{0}\in\mathbb{R}$ and any $z_{0}\in H^{1}_{0}(-1,1)$, see Proposition 2.7 in [1]. Moreover, there is a positive constant $C$ depending on $\overline{\psi},\overline{h},\mu$ and $T$ such that
$\|z\|_{H_{0}^{2,1}(\tilde{Q})}+\|k\|_{H^{1}(0,T)}\leq C\left(\|f_{0}\|_{L^{2}(\tilde{Q})}+\|v\|_{L^{2}(\omega\times(0,T))}+\|g_{0}\|_{L^{2}(0,T)}+|k_{0}|+\|z_{0}\|_{H^{1}_{0}(-1,1)}\right).$ (16)
2.2 Adjoint system of (14)
To determine the adjoint system of (14), multiply the first equation of (14) by $w:=w(x,t)$ and the third equation by $\theta:=\theta(t)$ and proceeding with integrations by parts, we find that the adjoint system of (14) is given by
$\left\{\begin{aligned} &-(\overline{h}w)_{t}-w_{xx}+\frac{(\overline{h})^{\prime}}{2}(xw)_{x}+nw=f_{1}&&\text{in}\;\tilde{Q},\\ &w(-1,t)=0&&\text{in}\;(0,T),\\ &w(1,t)=\mu\int_{-1}^{1}x\overline{\psi}_{x}(x,t)w(x,t)\mathrm{d}x+2\mu\theta(t)&&\text{in}\;(0,T),\\ &-\theta^{\prime}(t)+\int_{-1}^{1}m(x,t)w(x,t)\mathrm{d}x=g_{1}&&\text{in}\;(0,T),\\ &\theta(T)=\theta_{T},\\ &w(\cdot,T)=w_{T},&&\text{in}\;(-1,1).\end{aligned}\right$. (17)
@@ 2.2 Adjoint system of (14)
$w_{T}(-1)=0\quad\mbox{and}\quad w_{T}(1)=\mu\int_{-1}^{1}x\overline{\psi}_{x}(x,T)w_{T}(x)\mathrm{d}x+2\mu\theta_{T}.$ (18)
The following result assures the existence of strong solutions of (17) and its proof is a direct application of [1, Proposition 2.5] based on Leray-Schauder’s Fixed-Point Principle.
Proposition 3.
Let $w_{T}\in H^{1}(-1,1)$ and $\theta_{T}\in\mathbb{R}$ such that the compatibility condition (18) is satisfied and $(f_{1},g_{1})\in L^{2}(\tilde{Q})\times L^{2}(0,T)$, then (17) has a unique strong solution $(w,\theta)\in H^{2,1}(\tilde{Q})\times H^{1}(0,T)$, which satisfies the following estimate:
$\|w\|_{H^{2,1}(\tilde{Q})}+\|\theta\|_{H^{1}(0,T)}\leq C\left(\|f_{1}\|_{L^{2}(\tilde{Q})}+\|g_{1}\|_{L^{2}(0,T)}+|\theta_{T}|+\|w_{T}\|_{H^{1}(-1,1)}\right),$ (19)
@@
2.3 Null controllability of system (14)
The objective of this section is to prove null controllability for system (14) and some state estimation and control regularity that requires certain regularity of the initial data and inhomogeneous terms. We consider the following positive weight functions $\alpha$ and $\xi$, which depend on $\omega$:
@@ 2.3 Null controllability of system (14)
$\eta>0\text{ in }[-1,1],\quad\min_{x\in[-1,1]\backslash\omega^{\prime}}\left|\eta_{x}(x)\right|>0,\quad\eta(-1)=\eta(1)=\min_{x\in[-1,1]}\eta(x),$ (20)
even if we divide by $\eta||_{\infty}$, we can also assume that $\eta||_{\infty}=1$, where $\omega^{\prime}\subset\subset\omega$ is an open non empty subset. The existence of such a function $\eta$ satisfying (20) is proved in [20].
Notation 2.
@@
Proposition 4.
There exist constants $\lambda_{1},s_{1}\geq 1$ and $C_{1}>0$ such that, for any $\lambda\geq\lambda_{1}$, any $s\geq s_{1}(T+T^{2})$, any $w_{T}\in H^{1}(-1,1)$ and any $\theta_{T}\in\mathbb{R}$ with compatibility condition (18) is satisfied and any source terms $(f_{1},g_{1})\in L^{2}(\tilde{Q})\times L^{2}(0,T)$, the strong solution to (17) satisfies:
@@
$\leq C_{1}\left(s^{3}\lambda^{4}\iint_{\omega\times(0,T)}e^{-2s\alpha}\xi^{3}|w|^{2}\mathrm{d}x\mathrm{d}t+\iint_{\tilde{Q}}e^{-2s\alpha}|f_{1}|^{2}\mathrm{d}x\mathrm{d}t+\int_{0}^{T}e^{-2s\hat{\alpha}}|g_{1}|^{2}\mathrm{d}t\right).$ (21)
@@
System (17) can be written in the form
$\left\{\begin{aligned} &\varphi_{t}+d(t)\varphi_{xx}=f&&\text{in}\;\tilde{Q},\\ &\varphi(-1,t)=0&&\text{in}\;(0,T),\\ &\varphi(1,t)=\gamma(t)+\int_{-1}^{1}N(x,t)\varphi(x,t)\mathrm{d}x&&\text{in}\;(0,T),\\ &\gamma^{\prime}(t)-\int_{-1}^{1}R(x,t)\varphi(x,t)\mathrm{d}x=g&&\text{in}\;(0,T),\\ &\gamma(T)=2\mu\theta_{T},\\ &\varphi(\cdot,T)=\varphi_{T}&&\text{in}\;(-1,1),\end{aligned}\right$. (22)
where $(\varphi,\gamma):=(w,2\mu\theta)$, $d:=(\overline{h})^{-1},\;N:=\mu x\overline{\psi}_{x},\;R:=2\mu m$, $f:=-d\left(f_{1}+\left(\frac{(\overline{h})^{\prime}}{2}-n\right)\varphi-\frac{x}{2}(\overline{h})^{\prime}\varphi_{x}\right)$, $g:=-2\mu g_{1}$, $m$ and $n$ are defined in (9). Since $(\overline{\psi},\overline{h})$ is a smooth data (see (12)) such that $\overline{h}>L_{\star}^{2}$, using the Sobolev continuous embedding $H^{2,1}(\tilde{Q})\hookrightarrow C([0,T];H^{1}(-1,1))$ and $H^{1}(-1,1)$ is a Banach algebra, we obtain
@@
Note that the regularity $W^{1,\infty}(0,T)$ is sufficient for $d$. Hence, we are in a position to apply Theorem 3.1 of [1], so there exist constants $\lambda_{1},s_{1}\geq 1$ and $C>0$ such that, for any $\lambda\geq\lambda_{1}$, any $s\geq s_{1}(T+T^{2})$, the strong solution to (22) satisfies:
@@
We will deduce a Carleman estimate with functions blowing up only at $t=T$, which allows us to prove an observability inequality for system (17). Define the new weight functions:
@@
Note that $\ell\in C^{1}([0,T])$. An estimate with such weights is given in the following result. In the proof, we will use Proposition 4 and energy estimates (19).
Proposition 5.
There exist constants $\lambda_{1},s_{1}\geq 1$ such that for any $s\geq s_{1}$, any $\lambda\geq\lambda_{1}$, there exists a constant $C:=C(s,\lambda,T)$, which satisfies the following property: for any strong solution $(w,\theta)$ of (17), we have the following estimate:
@@
$\leq C\left(\iint_{\omega\times(0,T)}e^{-2s\beta}\ell^{-3}|w|^{2}\mathrm{d}x\mathrm{d}t+\iint_{\tilde{Q}}e^{-2s\beta}|f_{1}|^{2}\mathrm{d}x\mathrm{d}t+\int_{0}^{T}e^{-2s\hat{\beta}}|g_{1}|^{2}\mathrm{d}t\right).$ (23)
with $s_{1}$ and $\lambda_{1}$ as in Proposition 4 and $\hat{\beta}=\displaystyle\max_{x\in[-1,1]}\beta(x,\cdot)$.
@@
To do so, we split the left-hand side of (23) into two parts, corresponding to the restrictions of $(w,\theta)$ on $1,1]\times\left(0,T/2\right)$ and $1,1]\times\left(T/2,T\right)$.
Firstly, on $1,1]\times\left(T/2,T\right)$. Using Carleman’s estimate (21), $\alpha=\beta$ and $\xi=e^{\lambda(\kappa+\eta(x))}\ell^{-1}$ in $1,1]\times\left(T/2,T\right)$, we obtain
@@
$\leq C\left(\iint_{\omega\times(0,T)}e^{-2s\alpha}\xi^{3}|w|^{2}\mathrm{d}x\mathrm{d}t+\iint_{\tilde{Q}}e^{-2s\alpha}|f_{1}|^{2}\mathrm{d}x\mathrm{d}t+\int_{0}^{T}e^{-2s\hat{\alpha}}|g_{1}|^{2}\mathrm{d}t\right).$ (24)
@@
$\alpha\geq\beta\quad\mbox{in}\;\;[-1,1]\times(0,T)\quad\mbox{and}\quad\hat{\alpha}\geq\hat{\beta}\quad\mbox{in}\;\;(0,T).$ (25)
@@
$e^{-2s\alpha}\xi^{3}\leq Ce^{-2s\beta}\ell^{-3}\quad\mbox{in}\;\;[-1,1]\times(0,T),$ (26)
where $C>0$ only depending on $s$, $\lambda$ and $T$. Considering (24)-(26), we obtain
@@
Secondly, on $1,1]\times\left(0,T/2\right)$. Let us define the function $\vartheta\in C^{1}([0,T])$ such that
@@
$\left\{\begin{aligned} &-(\overline{h}w^{\vartheta})_{t}-w^{\vartheta}_{xx}+\frac{(\overline{h})^{\prime}}{2}(xw^{\vartheta})_{x}+nw^{\vartheta}=f^{\vartheta}&&\text{in}\;\tilde{Q},\\ &w^{\vartheta}(-1,t)=0&&\text{in}\;(0,T),\\ &w^{\vartheta}(1,t)=\mu\int_{-1}^{1}x\overline{\psi}_{x}(x,t)w^{\vartheta}(x,t)\mathrm{d}x+2\mu\theta^{\vartheta}(t)&&\text{in}\;(0,T),\\ &-\theta^{\vartheta}_{t}(t)+\int_{-1}^{1}m(x,t)w^{\vartheta}(x,t)\mathrm{d}x=g^{\vartheta}&&\text{in}\;(0,T),\\ &\theta^{\vartheta}(T)=0,\\ &w^{\vartheta}(\cdot,T)=0&&\text{in}\;(-1,1),\end{aligned}\right$.
where $(f^{\vartheta},g^{\vartheta}):=\vartheta(f_{1},g_{1})-\vartheta_{t}(\overline{h}w,\theta)\in L^{2}(\tilde{Q})\times L^{2}(0,T)$. From energy estimate (19), one has
@@
$\quad\left.+\|g_{1}\|_{L^{2}(0,3T/4)}+\|w\|_{L^{2}((-1,1)\times(T/2,3T/4))}+\|\theta\|_{L^{2}(T/2,3T/4)}\right).$ (27)
Using the boundedness from above and from below of the weight functions $\beta$, $\hat{\beta}$ and $\ell$ in $1,1]\times\left(0,3T/4\right)$ and (27), we obtain
@@
which, combined with (27), yields (23). ∎
@@
Proposition 6.
Let $(z_{0},k_{0})\in H^{1}_{0}(-1,1)\times\mathbb{R}$ and $(f_{0},g_{0})$ are source terms such that $\rho(f_{0},g_{0})\in L^{2}(\tilde{Q})\times L^{2}(0,T)$. Then, there exists a control $v^{*}$ such that the strong solution $(z^{*},k^{*})$ of (14) corresponding to $v^{*}$, $(f_{0},g_{0})$ and $(z_{0},k_{0})$, satisfies:
@@
$\quad\quad\leq C\left(\|\rho f_{0}\|_{L^{2}(\tilde{Q})}^{2}+\|\rho g_{0}\|_{L^{2}(0,T)}^{2}+\|z_{0}\|_{L^{2}(0,T)}^{2}+|k_{0}|^{2}\right).$ (28)
In particular (14) is null controllable. Moreover, $\rho_{2}(z^{*},k^{*})\in H^{2,1}(\tilde{Q})\times H^{1}(0,T)$
$\|\rho_{2}z^{*}\|^{2}_{H^{2,1}(\tilde{Q})}+\|\rho_{2}k^{*}\|^{2}_{H^{1}(0,T)}\leq C\left(\|\rho f_{0}\|_{L^{2}(\tilde{Q})}^{2}+\|\rho g_{0}\|_{L^{2}(0,T)}^{2}+\|z_{0}\|_{L^{2}(0,T)}^{2}+|k_{0}|^{2}\right).$ (29)